# In-situ Straining of Epitaxial Freestanding Ferroic Films through a MEMS Device

### v1

@@ Abstract
Mechanical strain can be used to control physical properties in materials. The experimental investigation of strain-induced effects at the nanoscale is of importance not only for its fundamental aspect, but also for the development of device applications. Transmission X-ray microscopy is a particularly well-suited technique for the nanoscale imaging of magnetic materials, but its compatibility with in-situ mechanical straining of samples is limited. In this work, we present a setup for applying tailored in-situ mechanical strains to freestanding thin films by means of a micro electromechanical system (MEMS) actuator. We then present a proof-of-concept experiment where a freestanding 80 nm thick (001) BiFeO3 multiferroic thin film is strained with the MEMS device, allowing us to control the coupled ferroelectric/spin cycloidal configuration.
@@ I Introduction
Examples of magnetoelastic effects whose investigation requires a microscopy technique include the study of domain wall and vortex dynamics as a function of applied strain [13, 14, 9], the study of strain-induced phase transitions in antiferromagnetic materials [23, 17], and the investigation of the the effect of strain on the coupling mechanisms in ferroic systems [18, 34, 20, 24, 33, 28]. Their investigation requires the combination of a non-invasive microscopy technique able to resolve nanometric features with contrast mechanisms that allow for the visualization of the desired ferroic order. Thanks to the possibility of combining non-invasive imaging with strong dichroic contrast mechanisms (X-ray magnetic circular and linear dichroism - XMCD/XLD), with the possibility of attaining nanometric resolutions, and with the option to apply in-situ excitations, soft X-ray microscopy techniques such as photoemission electron microscopy (PEEM) [7, 10, 9], scanning transmission X-ray microscopy (STXM) [13, 23, 17], and X-ray ptychography [6, 25] are an ideal match. In this work, we will concentrate on STXM and ptychography.
@@
STXM and X-ray ptychography are transmission microscopy techniques. Due to the low penetration depth of soft X-rays, only sample thicknesses well below the micrometer can be investigated. By using phase contrast imaging at photon energies below the elemental absorption edges of the sample, the thickness that can still be imaged with dichroic contrast can be extended to the micrometer range [25], but not significantly beyond that. Therefore, the standard method employed for straining a sample in non-transmission microscopy techniques, i.e. by means of a piezoelectric crystal actuator [10, 7, 14], cannot be used for transmission X-ray microscopy due to the non-negligible thickness of the piezoelectric crystal.
@@
Over the last years, a solution for in-situ mechanical straining of X-ray transparent samples was developed by modifying a gas cell used for atmospheric chemistry investigations [22]. With this method, a 50 nm thick Si3N4 membrane is bent by pressurizing the gas cell, as shown in the sketch in Fig. 1(a) [12]. Upon the application of a pressure difference between its two surfaces, the membrane buckles towards the region with lower pressure, generating a tensile strain at the top surface, and a compressive strain at the bottom surface, which is then transferred to structures on the membrane [12, 17]. Due to the fabrication constraints, the typical geometries of Si3N4 membranes are square or rectangular, i.e. the strain is applied primarily along the two axes of the square/rectangle [12, 13, 17]. By tailoring the membrane geometry, it is possible e.g. to simulate an effective uniaxial strain application [12, 17], or to simulate a situation as close as possible to a circular membrane, resulting in the generation of an isotropic strain [12].
Figure 1: (a) Sketch of the gas cell-based setup used to generate strain through the bending of a Si3N4 membrane. Image adapted from [12]. (b) XMCD-STXM images of a Co40Fe40B20 microstructured square fabricated on a rectangular Si3N4 membrane under different pressure differences between the two sides of the membrane (measured by gauges G1 and G2 in (a)), generating a uniaxial tensile strain of different magnitudes. The red arrows mark the direction of the in-plane magnetization in the microstructured squares. Image adapted from [13].
@@
Membrane bending has been used for the uniaxial straining of Co40Fe40B20 microstructured elements stabilizing a Landau flux-closure state [12]. Using a rectangular membrane, a uniaxial magnetic anisotropy due to the magnetoelastic effect could be generated, as shown in Fig. 1(b) [12, 13]. Another example is the demonstration of strain-driven control of the antiferromagnetic configuration in freestanding $\alpha$-Fe2O3 thin films [23, 17].
@@
While STXM investigations combined with bending of Si3N4 membranes were possible [13, 17], this setup has limitations that reduce its applicability to epitaxial systems. The main limitation is that the Si3N4 membranes are amorphous and can therefore not be used as substrates for the growth of epitaxial films. While single-crystalline Si membranes are commercially available, the growth of epitaxial materials on these is challenging due to the usual requirement of heating the substrate to high temperatures during the growth, and the range of materials that can be epitaxially grown on Si is limited, e.g. excluding many ferroic perovskites. In addition, the Si3N4 membrane geometries are restrictive: while certain complex geometries such as e.g. circular membranes are feasible [29], it is not possible to fabricate membranes with arbitrary geometries, limiting the ensemble of strain configurations that can be generated.
@@
In this work, we present a different approach to the in-situ straining of X-ray transparent samples, based on the use of micro electromechanical system (MEMS) devices on which a lamella was attached by means of ion-beam assisted deposition of carbon layers in a focused ion beam (FIB) tool, in order to fix the two ends of the lamella on the MEMS device. By electrically poling the MEMS device, the X-ray transparent lamella can be strained to values significantly above those achievable either by piezoelectric crystal actuators or by membrane bending. Below, we present the device fabrication and the proof-of-concept experiment demonstrating the application of a tensile strain onto a freestanding BiFeO3 (BFO) lamella, leading to strain-induced motion of ferroelectric domain walls and changes to the coupled spin cycloid state.
@@ II In-situ straining with a MEMS device
For the proof-of-concept experiments described in this work, a BFO lamella extracted from a 80 nm thick freestanding (001) BFO film (details on the BFO film growth and release from the substrate can be found in [6]) needed to be transferred onto the gap between the two cantilevers. For this preparation, a Ga-FIB instrument (Zeiss NVision40) equipped with a micromanipulator from Kleindieck was employed.
As the freestanding films are deposited onto a standard 3 mm diameter Cu mesh after their release from the substrate in a potentially random orientation, it is necessary to first determine the orientation of the film, in order to guarantee that the lamella is cut along the desired [1-10]/[110] directions. To achieve this, the freestanding film was measured in advance by means of XLD ptychography to identify the crystallographic directions of the BFO film from the orientation of the spin cycloid and from the orientation of the edge of the film [6, 4]. During the extraction of the lamella, minimal damage from the ion beam has been achieved by using a 10 pA/30 kV beam current for all steps where the membrane had to be imaged with the ion beam. Two line cuts at a distance of 10-20 µm and a length of 80-100 µm were performed with a focused Ga-ion beam set to a 300 pA/30 kV beam current. The micromanipulator needle was then attached with ion-assisted carbon deposition close to the middle of the lamella, which helped to reduce the bending of the very thin stripe after cutting its end off from the rest of the membrane. After cutting the lamella free and lifting it out from the TEM-grid, it was fixed with ion assisted deposition of carbon at its destination, i.e. on both sides of the gap on the MEMS device, followed by cutting off the manipulator needle. Having the manipulator attached in the middle of the stripe had the additional advantage of helping to not damage the MEMS cantilever structures upon cutting off the manipulator needle from the stripe. A scanning electron micrograph (SEM) of the transferred lamella is shown in Fig. 3.
@@
The images were then reconstructed from the recorded diffraction patterns using the PtychoShelves Matlab package [32], using combined difference mapping and maximum likelihood algorithms [31]. In order to account for the incomplete coherence of the X-ray beam, three probe modes were utilized. The reconstruction yields a complex-valued image, of which the amplitude contrast was selected for the ptychographic XLD images presented here. An example of a reconstructed image of the BFO lamella is shown in Fig. 4(a), with the XLD contrast variations caused by the spin cycloid shown in Fig. 4(b).
@@
Figure 5: (a) Variation of the distance between the two MEMS cantilevers as a function of the applied voltage difference between the cantilevers’ top and bottom surfaces. The recorded displacement is independent of the size of the gap between the two cantilevers; (b) Maximum tensile strain generated by the displacement of the cantilevers.
The MEMS devices were poled with voltages from 0 to 40 V in steps of 5 V. For each voltage step, a SEM image of each cantilever was acquired (images available in the Supplementary Information). The voltage-dependent displacement of the cantilevers was determined from the images and is shown in Fig. 5(a). Three different cantilevers, marked as devices 1-3 in the figure, with two differently sized gaps, were investigated. As it is possible to observe in Fig. 5(a), the displacement does not depend on the size of the cut, but only on the applied voltage. The size of the cut determines the maximum strain that can be achieved, as shown in Fig. 5(b), where the applied strain is defined as $\Delta\varepsilon/\varepsilon_{0}$, being $\Delta\varepsilon$ the displacement, and $\varepsilon_{0}$ the gap size at no applied voltage. As device 2 has a smaller gap (about 15 µm, against about 25 µm for the other two devices), the achievable strain is higher than for the other two cantilever devices. All of the MEMS devices shown in Fig. 5 can generate significantly larger tensile strains than both single-crystal piezoelectric substrates and membrane bending geometries.
The tensile strain generated by the MEMS device follows a parabolic dependence with respect to the applied voltage (see Fig. 5). If the same device is used for hundreds of cycles, the strain-vs-voltage response must exhibit a more linear behavior that needs to be taken into consideration if aging experiments are performed or if the same sample is investigated in several measurement campaigns [26].
III Rotation of ferroelectric domain walls through tensile strain
The image shown in Fig. 4 was acquired in the as-transferred configuration, i.e. without any applied strain. Upon poling the MEMS device, a tensile strain of varying magnitudes was applied to the BFO lamella. At each strain level, gauged as described above from the displacement of the cantilevers, an XLD-ptychography image of the ferroelectric and spin cycloidal configuration of the BFO lamella was acquired. While no global changes in either the orientation or the period of the spin cycloid of the BFO lamella were observed, there exist several regions where the mechanical straining leads to ferroelectric domain wall motion.
@@ III Rotation of ferroelectric domain walls through tensile strain
Fig. 6 shows one of the regions where ferroelectric domain wall motion was observed. For this domain wall, the motion occurred upon the application of a tensile strain of about 1%, as shown in Fig. 6(a). The two ferroelectric domains separated by the moving domain wall exhibit both a rotation of the spin cycloid propagation vector and a reduction of the cycloid period, likely indicating a rotation of the coupled ferroelectric polarization direction either towards the strain axis or perpendicular to it. Similar changes occur for other ferroelectric domains in the imaged ROI (see Supplementary Information).
@@ IV Conclusion and Outlooks
In conclusion, we have shown a design for a MEMS device that can be used for the soft X-ray ptychographic imaging of in-situ strained epitaxial freestanding films. The concept is based on a dual-cantilever MEMS device with a 15-25 µm gap between the cantilevers, where an epitaxial film lamella is positioned by means of a FIB instrument. By poling the MEMS device, a tensile strain in the axis parallel to the cantilevers can be transferred to the lamella, whilst preserving the X-ray transparency requirement for ptychography imaging. Tensile strains reaching several percentage points were obtained, i.e. above those achievable with other methods compatible with transmission microscopy. A proof-of-concept measurement was performed with a 80 nm thick (001) BFO lamella, where the application of a tensile strain of ca. 1% led to the combined motion of several ferroelectric domain walls, leading also to a change in the orientation and period of the spin cycloid coupled with the ferroelectric order in the BFO.
The design shown above allows one to apply a tensile strain to lamellas. However, other types of strain such as compressive or shear strain are also of interest. To generate shear strain, the same MEMS device geometry could be exploited by actuating the piezoelectric patches with differential voltage as sketched in Fig. 7. This could induce a rotation of the two cantilevers and a consequential torque on the film under test. Additionally, a direct gauging of the applied strain could be implemented by measuring the displacement of the cantilever with the piezoresistive sensor integrated with the actuator, as shown for the MEMS micromirrors reported in [15]. Another possible improvement is in the use of a piezoelectric material such as AlN, which could ensure a more linear response of the system, thus reducing the need for an in-situ gauging of the strain by means of focus or linescans whilst also enabling bipolar actuation.
@@
Finally, another possible advantage in using MEMS actuators for the in-situ straining of materials is that their mechanical resonances can be controlled: while the resonance frequency for the device shown in this work is relatively low (the first bending mode is around 10 kHz), MEMS devices can be designed with resonance frequencies up to the MHz range [1]. Such designs, if modified to be compatible with transmission microscopy, could enable the time-resolved investigation of magnetoelastic processes, both with standard STXM time-resolved imaging [27, 11] and, if the resonance frequency is tuned to one of the harmonics of the master clock of the synchrotron lightsource, time-resolved ptychography [3].

### v2

# In situ straining of epitaxial freestanding ferroic films by a microelectromechanical device

@@ Abstract
Mechanical strain can be used to control physical properties in materials. The experimental investigation of strain-induced effects at the nanoscale is of importance not only for its fundamental aspects, but also for the development of device applications. Transmission X-ray microscopy is a particularly well-suited technique for nanoscale imaging of magnetic materials, but its compatibility with in-situ mechanical straining of samples is limited. In this work, we present a setup for applying tailored in-situ mechanical strains to freestanding thin films by means of a micro electromechanical system (MEMS) actuator. We then present a proof-of-concept experiment in which a freestanding 80 nm thick (001) BiFeO3 multiferroic thin film is strained with the MEMS device, allowing us to control the coupled ferroelectric/spin cycloidal configuration.
@@ I Introduction
Examples of magnetoelastic effects, the investigation of which requires a microscopy technique include the study of domain wall and vortex dynamics as a function of applied strain [3, 4, 5], the study of strain-induced phase transitions in antiferromagnetic materials [6, 7], and the investigation of the effect of strain on the coupling mechanisms in ferroic systems [8, 9, 10, 11, 12, 13]. Their investigation requires the combination of a non-invasive microscopy technique able to resolve nanometric features with contrast mechanisms that allow for the visualization of the desired ferroic order. Thanks to the possibility of combining non-invasive imaging with strong dichroic contrast mechanisms (X-ray magnetic circular or X-ray linear dichroism - XMCD/XLD), with the possibility of attaining nanometric resolutions, and with the option to apply in-situ excitations, soft X-ray microscopy techniques such as photoemission electron microscopy (PEEM) [14, 15, 5], scanning transmission X-ray microscopy (STXM) [3, 6, 7], and X-ray ptychography [16, 17] are an ideal match. In this work, we will concentrate on STXM and ptychography.
@@
STXM and X-ray ptychography are transmission microscopy techniques. Due to the low penetration depth of soft X-rays, only sample thicknesses under one micrometer can be investigated. Using phase contrast imaging at photon energies below the elemental absorption edges of the sample, the thickness that can still be imaged with dichroic contrast can be extended to the micrometer range [17], but not significantly beyond that. Therefore, the standard method employed for straining samples in non-transmission microscopy techniques, i.e. by means of a piezoelectric crystal actuator [15, 14, 4], cannot be used for transmission X-ray microscopy due to the non-negligible thickness of the piezoelectric crystal.
@@
In recent years, a solution for in-situ mechanical straining of X-ray transparent samples was developed by modifying a gas cell used for atmospheric chemistry investigations [18]. With this method, a 50 nm thick Si3N4 membrane is bent by pressurizing the gas cell, which buckles towards the region with lower pressure, generating a tensile strain at the top surface, and a compressive strain at the bottom surface, which is then transferred to structures on the membrane [19, 7]. Due to the fabrication constraints, the typical geometries of Si3N4 membranes are square or rectangular, i.e. the strain is applied primarily along the two axes of the square/rectangle [19, 3, 7]. By tailoring the membrane geometry, it is possible e.g. to simulate an effective uniaxial strain application [19, 7], or to simulate a situation as close as possible to a circular membrane, resulting in the generation of an isotropic strain [19].
@@
Membrane bending has been used for the uniaxial straining of Co40Fe40B20 microstructured elements stabilizing a Landau flux-closure state [19]. Using a rectangular membrane, a uniaxial magnetic anisotropy due to the magnetoelastic effect could be generated [19, 3]. Another example is the demonstration of strain-driven control of the antiferromagnetic configuration in freestanding $\alpha$-Fe2O3 thin films [6, 7].
@@
While STXM investigations combined with bending of Si3N4 membranes were possible [3, 7], this setup has limitations that reduce its applicability to epitaxial systems. The main limitation is that the Si3N4 membranes are amorphous and can therefore not be used as substrates for the growth of epitaxial films. While single-crystalline Si membranes are commercially available and two dimensional nanosheets used as growth template layer can allow for the growth of some epitaxial material [20, 21], this still holds several experimental challenges. In addition, the Si3N4 membrane geometries are restrictive: while certain complex geometries such as e.g. circular membranes are feasible [22], it is not possible to fabricate membranes with arbitrary geometries, limiting the ensemble of strain configurations that can be generated.
@@
As no commercial solutions for the in-situ straining of thin membranes in conditions compatible with soft X-ray microscopy exist, in this work we present a different approach to the in-situ straining of X-ray transparent samples, based on the use of micro electromechanical system (MEMS) devices. A lamella was attached to these by means of ion-beam assisted deposition of carbon layers in a focused ion beam (FIB) tool, fixing the two ends of the lamella on the MEMS device. By electrically poling the MEMS device, the X-ray transparent lamella can be strained to values significantly above those achievable either by piezoelectric crystal actuators or by membrane bending. Below, we present the device fabrication and the proof-of-concept experiment demonstrating the application of a tensile strain onto a freestanding BiFeO3 (BFO) lamella, leading to strain-induced motion of ferroelectric domain walls and changes to the coupled spin cycloid state.
@@ II In-situ straining with a MEMS device
For the proof-of-concept experiments described in this work, a lamella extracted from a 80 nm thick freestanding (001) BFO film needed to be transferred onto the gap between the two cantilevers. Epitaxial BFO/Sr3Al2O6 (SAO) heterostructures were grown on (001) oriented SrTiO3 (STO) substrates using a pulsed laser deposition system (Mobile Combi-Laser MBE MC-LMBE, Pascal Co., Ltd.). A KrF excimer laser ($\lambda$ = 248 nm) with an energy of 200 mJ and a repetition rate of 10 Hz was utilized for all growth processes. First, a 25 nm sacrificial layer of SAO was deposited on the STO substrate at 740 °C under an oxygen pressure of 45 mTorr. Subsequently, an 80 nm BFO layer was deposited on top of the SAO layer at 650 °C under an oxygen pressure of 65 mTorr. Following the growth, the sample was annealed at 400 C in nearly 1 atm of oxygen for 20 minutes to reduce oxygen vacancies, followed by a slow cooling process to room temperature under the same oxygen environment. Finally, the sample was immersed in deionized water to dissolve the sacrificial SAO layer, allowing the freestanding BFO film to be released and transferred from the solution onto a TEM grid with a lacy carbon support for subsequent XLD analysis.
For the transfer of the BFO freestanding films, a Ga-FIB instrument (Zeiss NVision40) equipped with a micromanipulator from Kleindieck was employed. As the freestanding films are deposited onto a standard 3 mm diameter Cu mesh after their release from the substrate in a potentially random in-plane orientation, it is necessary to first determine the orientation of the film, in order to guarantee that the lamella is cut along the desired [1-10]/[110] directions. To achieve this, the freestanding film was measured in advance by means of XLD ptychography to identify the crystallographic directions of the BFO film from the orientation of the spin cycloid and from the orientation of the edge of the film [16, 25]. During the extraction of the lamella, minimal damage from the ion beam has been achieved by using a 10 pA/30 kV beam current for all steps where the membrane had to be imaged with the ion beam. Two line cuts at a distance of 10-20 µm and a length of 80-100 µm were performed with a focused Ga-ion beam set to a 300 pA/30 kV beam current. The micromanipulator needle was then attached with ion-assisted carbon deposition close to the middle of the lamella, which helped to reduce the bending of the very thin stripe after cutting its end off from the rest of the membrane. After cutting the lamella free and lifting it from the TEM-grid, it was fixed with ion-assisted deposition of carbon on both sides of the gap on the MEMS device, followed by cutting off the manipulator needle. The attachment of the manipulator to the middle of the stripe had the additional advantage of helping to not damage the MEMS cantilever structures upon cutting off the manipulator needle from the stripe. A scanning electron micrograph (SEM) of the transferred lamella is shown in Fig. 2.
@@
The images were then reconstructed from the recorded diffraction patterns using the PtychoShelves software package [30] with the combined difference map [31] and maximum likelihood algorithms [32]. In order to account for the incomplete coherence of the X-ray beam, three probe modes were utilized [33]. The ptychographic reconstruction yielded a complex-valued image, of which the amplitude contrast was selected for the ptychographic XLD images presented here. An example of a reconstructed image of the BFO lamella is shown in Fig. 3(a), with the XLD contrast variations caused by the spin cycloid shown in Fig. 3(b).
@@
Figure 4: (a) Variation of the distance between the two MEMS cantilevers as a function of the applied voltage difference between the top and bottom surfaces of the cantilever determined from SEM images as a function of the applied voltage. The recorded displacement is independent of the size of the gap between the two cantilevers; (b) Maximum tensile strain generated by the displacement of the cantilevers.
The MEMS devices were poled with voltages from 0 to 40 V in steps of 5 V. For each voltage step, an SEM image of each cantilever was acquired (images in the Supplementary Information [34]). The voltage-dependent displacement of the cantilevers was determined from the images and is shown in Fig. 4(a). Three different cantilevers, marked as devices 1–3 in the figure, with two differently sized gaps, were investigated. As can be observed in Fig. 4(a), the displacement does not depend on the size of the cut, but only on the applied voltage. Note that the errors in the gap distance are due to the pixel size in the SEM images. The size of the cut determines the maximum strain that can be achieved, as shown in Fig. 4(b), where the applied tensile strain is defined as $\Delta x/x_{0}$, being $\Delta x$ the displacement, and $x_{0}$ the gap size at no applied voltage. As device 2 has a smaller gap (about 15 µm, against about 25 µm for the other two devices), the achievable strain is higher than for the other two cantilever devices. All of the MEMS devices shown in Fig. 4 can generate significantly larger tensile strains than both single-crystal piezoelectric substrates and membrane bending geometries and comparable to the current state-of-the-art for perovskite films, which is about 8% [35].
The tensile strain generated by the MEMS device follows a parabolic dependence with respect to the applied voltage (see Fig. 4). If the same device is used for hundreds of cycles, the strain-vs-voltage response changes with respect to the first cycles. This change, typically driven by wake-up/conditioning effects linked to the reorganization of the ferroelectric domain walls and defects in the piezoelectric material of the MEMS, lead to a linearization of the strain-vs-voltage response, which needs to be taken into consideration if the same sample is investigated in several measurement campaigns [24, 36].
III Rotation of ferroelectric domain walls by tensile strain
The image shown in Fig. 3 was acquired in the as-transferred configuration, i.e. without any applied strain. Upon poling the MEMS device, a tensile strain of varying magnitudes was applied to the BFO lamella. At each strain level, gauged as described above from the displacement of the cantilevers, an XLD-ptychography image of the ferroelectric and spin cycloidal configuration of the BFO lamella was acquired. While no global changes in either the orientation or the period of the spin cycloid of the BFO lamella were observed, several regions existed where the mechanical straining led to ferroelectric domain wall motion.
@@ III Rotation of ferroelectric domain walls by tensile strain
Fig. 5 shows one of the regions where ferroelectric domain wall motion was observed. For this domain wall, the motion occurred upon the application of a tensile strain of about 1%, as shown in Fig. 5(a). The two ferroelectric domains separated by the moving domain wall exhibit both a rotation of the spin cycloid propagation vector and a reduction of the cycloid period, likely indicating a rotation of the coupled ferroelectric polarization direction either towards the strain axis or perpendicular to it. Similar changes occur for other ferroelectric domains in the imaged ROI (see Supplementary Information [37]). A change in the cycloid period was also observed for substrate-bound BFO films when increasing the tensile strain component if compared to the growth conditions on a SrTiO3 substrate [8], indicating that similarities in the behavior of the spin cycloid with increasing tensile strain can be found between substrate-bound and freestanding films.
@@ IV Conclusion and Outlooks
In conclusion, we have shown a design for a MEMS device that can be used for the soft X-ray ptychographic imaging of in-situ strained epitaxial freestanding films. The concept is based on a dual-cantilever MEMS device with a 15–25 µm gap between the cantilevers, where an epitaxial film lamella is positioned by means of a FIB instrument. By poling the MEMS device, a tensile strain in the axis parallel to the cantilevers can be transferred to the lamella, whilst preserving the X-ray transparency requirement for ptychography imaging. Tensile strains reaching several percentage points were obtained, i.e. above those achievable with other methods compatible with transmission microscopy. A proof-of-concept measurement was performed with a 80 nm thick (001) BFO lamella, where the application of a tensile strain of ca. 1% led to the combined motion of several ferroelectric domain walls, also causing a change in the orientation and period of the spin cycloid coupled with the ferroelectric order in the BFO.
The design shown above allows one to apply a tensile strain to lamellas. However, other types of strain such as compressive or shear strain are also of interest. To generate shear strain, the same MEMS device geometry could be exploited by actuating the piezoelectric patches with differential voltage as sketched in Fig. 6. This could induce a rotation of the two cantilevers and a subsequent torque on the film under test. Additionally, a direct gauging of the applied strain could be implemented by measuring the displacement of the cantilever with the piezoresistive sensor integrated with the actuator, as shown in [38, 39, 40]. This would allow one not only to directly gauge the applied strain in-situ, but also to significantly reduce the error in the determined strain if compared to post-mortem SEM investigations. Another possible improvement is in the use of a piezoelectric material such as AlN, which could ensure a more linear response of the system, thus reducing the need for an in-situ gauging of the strain by means of focus or linescans whilst also enabling bipolar actuation.
@@
Finally, another possible advantage of using MEMS actuators for the in-situ straining of materials is that their mechanical resonances can be controlled: while the resonance frequency for the device shown in this work is relatively low (the first bending mode is around 10 kHz), MEMS devices can be designed with resonance frequencies up to the MHz range [41], allowing us to explore dynamical processes at MHz frequencies, such as e.g., the dynamics of the magnetoelastic effect [4]. Such designs, if modified to be compatible with transmission microscopy, would allow us to access the MHz frequency range and perform time-resolved investigation of magnetoelastic processes, both with standard STXM time-resolved imaging [42, 43] and, if the resonance frequency is tuned to one of the harmonics of the master clock of the synchrotron lightsource, time-resolved ptychography [44].